OpenStudy (anonymous):

Integral calculus: Find the area of the region enclosed by the cardiod r=20-20cos(theta) help appreciated :)

OpenStudy (anonymous):

Take the integral with respect to theta from 0 to 2pi?

OpenStudy (anonymous):

:) thanks, what about are the limits for r?

OpenStudy (blockcolder):

That's not it. Isn't the formula for area: \[\int_{\alpha}^{\beta}{1 \over 2}r^2 \text{d}\theta\] \[r^2=(20-20\cos\theta)^2=[20(1-\cos\theta)]^2=400(1-\cos\theta)^2\] \[A=200\int_0^{2\pi}1-2\cos\theta-\cos^2\theta \text{d}\theta\]

OpenStudy (anonymous):

i think this question requires \[\int\limits_{\alpha}^{\beta}\int\limits_{r1(\theta)}^{r2(\theta)} f(r,\theta) r drd \theta\] though....

OpenStudy (anonymous):

Eek.. My mistake.. Proper formula is actually the A = 1/2 r^2 theta.

OpenStudy (anonymous):

yeah seems logical blockcolder....i get very confused with polar form!!...j

OpenStudy (blockcolder):

Can you find the area now? :D

OpenStudy (anonymous):

Something tells me I need to put '1' as the function inside the integral...to find area. Im abit lost!

OpenStudy (anonymous):

My lecturers examples of this topic are a little unhelpful :(

OpenStudy (blockcolder):

Yes, the function in the double integral is indeed 1.

OpenStudy (anonymous):

Can I find the area of the half cardiod and use symmetry?

OpenStudy (blockcolder):

You can. :D

OpenStudy (anonymous):

So what would be my limits then (this is where i come unstuck) for theta and r?

OpenStudy (blockcolder):

For r, the limits are 0 and 20-20cos(theta). For theta, 0 and pi.

OpenStudy (anonymous):

okay and then I just muliply by 2..and walah we have area?

OpenStudy (blockcolder):

Yes. :D

OpenStudy (anonymous):

Thank you very much for the help guys, ill give it a go now!

OpenStudy (anonymous):

To make up for my bad answer in the first place.. I will give you this example from my textbook. Hope this helps: https://skitch.com/jrcdude/8w78n/stewart-calculus-early-transcendentals-7th-txtbk.pdf-page-1-032-of-1-356

OpenStudy (anonymous):

Haha, thanks Jason- appreciate it!

OpenStudy (anonymous):

Took me a little long but I got \[600pi \]

OpenStudy (anonymous):

Found a brilliant example for those interested :)

OpenStudy (anonymous):

and forgot to paste the link! haha. here it is! http://www.maths.manchester.ac.uk/service/MATH20003/cardioid.pdf